how many gal. of a 40% sugar solution must be mixed with 8 gal. of a 30% sugar solution to make a 32% solution

Answer by richwmiller(17219)

0.4*x+0.3*8=0.32(8+x)
0.4x-0.32x=2.56-2.4
0.08x=0.16
x=2
check
0.4*2+0.3*8=0.32(8+2)
0.8+2.4=0.32(10)
3.2=3.2
ok